POOLTYPE(1) POOLTYPE(1) NAME pooltype−displayaWEBpoolfile SYNOPSIS pooltypepoolfilename DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Thepooltypeprogramtranslatesa(program-oriented)WEBstringpoolfiletoa(human-ori-ented)textfile. Thepoolfilenamemustbecomplete;noaddingofdefaultextensionsorpathsearchingisdone. SEEALSO tangle(1). AUTHORS DonaldE.KnuthwrotetheprogramandoriginallyportedittoUnix.Itwaspublishedaspartof theTEXwaretechnicalreport,availablefromtheTEXUsersGroup. Web2C7.32x 27December1992 1